# Geometry and cohomology of compactified Deligne–Lusztig varieties

### v2

@@ 2.3. Criterion for connectedness
The $F$-orbit of $w\in W$ is the set ${{\rm Orb}_{F}(w)}:=\{F^{n}(w)|n\in\mathbb{Z}_{\geq 0}\}$. Under the assumption that $F(S)=S$, we define the set
@@
Let $w_{1},\ldots,w_{r}\in W$ and denote $I={{\text{supp}}}_{F}(w_{1},\ldots,w_{r})$. The Deligne–Lusztig varieties $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are irreducible if and only if $I=S_{F}$. In general, they have $|G^{F}/P_{I}^{F}|$ connected components, where $W_{I}$ is the $F$-stable parabolic subgroup of $W$ generated by $\{s\in W|{{\rm Orb}_{F}(s)}\in I\}$ and $P_{I}={B}W_{I}{B}$.
@@
via projection, cf. [DMR, §2.3.1]. It also follows that $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are dense open subschemes of ${\overline{X}^{S^{*}}(v)}$. Thus it suffices to prove that ${\overline{X}^{S^{*}}(v)}$ is connected if and only if $v$ has full $F$-support.
The proof showing that ${\overline{X}^{S^{*}}(v)}$ is connected is analogous to the proof of [BonnafeRouquierOntheirred] for classical Deligne–Lusztig varieties. Suppose $I=S_{F}$ and thus ${{\text{supp}}}_{F}(v)=S_{F}$. Let $Z$ be the connected component of ${\overline{X}^{S^{*}}(v)}$ which contains the point $({B},\ldots,{B})$. Denote the stabiliser of $Z$ in $G^{F}$ by ${\rm Stab}(Z)$. Fix $v^{\prime}$ to be of minimal length such that $v^{\prime}\preceq v$ and ${{\text{supp}}}_{F}(v^{\prime})=S_{F}$. In particular, $v^{\prime}$ is a reduced expression of $\alpha(v^{\prime})\in W$, which is a Coxeter element. This implies that ${\overline{X}^{S^{*}}(v^{\prime})}$ is irreducible [LusztigCoxeterOrbits, Proposition 4.8].
On the other hand, there is a $G^{F}$-equivariant embedding ${\overline{X}^{S^{*}}(v^{\prime})}\to{\overline{X}^{S^{*}}(v)}$. Let $Z^{\prime}$ be the connected component of ${\overline{X}^{S^{*}}(v^{\prime})}$ containing $({B},\ldots,{B})$. Then ${\rm Stab}(Z^{\prime})\subseteq{\rm Stab}(Z)$. However, ${\overline{X}^{S^{*}}(v^{\prime})}$ is irreducible, so ${\rm Stab}(Z^{\prime})=G^{F}$. Thus ${\overline{X}^{S^{*}}(v)}$ is irreducible. In this case, its dense open subschemes $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are also irreducible [GoerWedhAlgGeom1, Proposition 3.24].
Suppose ${{\text{supp}}}_{F}(w_{1},\ldots,w_{r})=I$ and so ${{\text{supp}}}_{F}(v)=I$. By [DMR, Proposition 2.3.8], we know that there is a $G^{F}$-equivariant isomorphism
@@
whose fibres are isomorphic to $\overline{X}^{S^{*}}_{L_{I}}(v)$. In this case, we know that $v$ has full $F$-support with respect to the Weyl group of $L_{I}$, so ${\overline{X}^{S^{*}}(v)}$ has $|G^{F}/P_{I}^{F}|$ connected components. ∎
@@ 2.4. $F_{p}$-regularity
In [Wang22, Proposition A.6], we showed that some compactified Deligne–Lusztig varieties are strongly $F_{p}$-regular. This property generalises to arbitrary compactified Deligne–Lusztig varieties. As in loc. cit., we use a similar strategy as in the proof of [Lusztigunipotentch, Lemma 4.3], which uses faithfully flat descent of the smoothness property along the Lang map.
@@
When $\overline{Bw_{i}B}$ is smooth for all $i$, this is proved by faithfully flat descent of smoothness in [Lusztigunipotentch, Lemma 4.3] and [DMR, Proposition 2.3.5], cf. [GoerWedhAlgGeom1, Proposition 14.59]. Analogously, we use faithfully flat descent of strongly $F_{p}$-regularity to show that any compactified Deligne–Lusztig variety is strongly $F_{p}$-regular.
For each $i=1,\ldots,r$, we know that $\overline{{B}w_{i}{B}}$ is strongly $F_{p}$-regular [BrTh, Cor. 4.1, p. 958] and thus Cohen-Macaulay [BrTh, Cor. 4.2, p. 959]. The product
@@
is strongly $F_{p}$-regular by faithfully flat descent along the projection map ${\rm pr}_{i}:\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}\to\overline{{B}w_{i}{B}}$ [HoHu, Theorem 5.5.b]. Recall that the Lang map $\mathcal{L}_{1}$ is étale, so its closed fibres are smooth and thus strongly $F_{p}$-regular, cf. [KSmith, Ch 3, Remark 1.6, Theorem 4.4(a)]. We may now apply [Aber, Theorem 3.6] and conclude that
@@
is strongly $F_{p}$-regular. Finally, by Lemma 2.2.1 and the faithfully flat descent of strong $F_{p}$-regularity [HoHu, Theorem 5.5], $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is strongly $F_{p}$-regular.
Strongly $F_{p}$-regular rings are $F_{p}$-rational, so $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is normal and Cohen-Macaulay by [HoHu, Theorem 4.2, 6.27]. ∎
@@
For $w\in W$, the Zariski closure $\overline{X(w)}$ in $G/B$ is not always smooth, but for any $v\in S^{*}$ that gives a reduced expression of $w$, the projection map ${\overline{X}^{S^{*}}(v)}\to\overline{X(w)}$ is a proper birational morphism [DL, Lemma 9.11].
@@ 3.1. Smoothness via braid relations
We consider some specific elements in the Weyl group for which the compactified Deligne–Lusztig variety is smooth. To begin, we recall the relations in the Weyl group of finite reductive groups, cf. [GeckPfeiffer].
All the relations in the group presentation of the Weyl group as a finite Coxeter group $(W,S)$ can be written as
$(st)^{m_{st}}=1,\ s,t\in S,$ (1)
where $m_{st}\in\mathbb{Z}$ such that $m_{ss}=1$ and $m_{st}\geq 2$ if $s\neq t$. In the case of $s\neq t$, we call (1) a braid relation in the Weyl group $W$. We refer to such an equation (1) as a braid relation in $W$.
Remark 3.1.1.
If $F(S)=S$, then whenever $(st)^{m_{st}}=1,\ s,t\in S$ is a braid relation in $W$, the equations
$(ts)^{m_{st}}=1,\hskip 14.22636pt\ (F(s)F(t))^{m_{st}}=1,\hskip 14.22636pt\text{and}\hskip 14.22636pt(F(t)F(s))^{m_{st}}=1$
also hold.
For $w=w_{1}w_{2}\in S^{*}$, we call $w_{1}$ a prefix of $w$ and $w_{2}$ a suffix of $w$.
Definition 3.1.2.
Let $(st)^{m}=1$, with $\ s,t\in S$ and $s\neq t$, be a braid relation in $W$. We define
$P(s,t):=st\cdots\in S^{*},$ (2)
such that ${{\ell^{S^{*}}}(P(s,t))}=m$ and $P(s,t)$ is a prefix of $(st)^{m}$ in $S^{*}$. When there is no ambiguity, we also use $P(s,t)$ to denote its canonical projection $\alpha(P(s,t))\in W$.
@@
Let $w_{1},\ldots,w_{r}\in W$. Suppose that for all $j\in\{1,\ldots,r\}$, we have $w_{j}=P(s,t)$ for some braid relation in $W$. Then the compactified Deligne–Lusztig variety $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth.
@@
By Proposition 2.4.1, it suffices to show that for all $j=1,\ldots,r$, the associated shceme $\overline{{B}w_{j}{B}}$ is smooth.
Consider $w_{j}=P(s,t)$ with $(st)^{m}=1$. Denote $J=\{s,t\}\subseteq S$. The the corresponding subgroup $W_{J}\subseteq W$ generated by $J$ has a group presentation
@@
Let $P_{J}=BW_{J}B\subseteq G$ be the parabolic subgroup corresponding to $W_{J}$. Since $w_{j}$ is the longest element in $W_{J}$, we know that $\overline{{B}w_{j}{B}}=P_{J}$ and is smooth cf. [DigneMichelBook, Proposition 3.2.8]. ∎
@@ 3.2. Higher direct images of the structure sheaf for birational morphisms
Definition 3.2.1.
For any $v\in S^{*}$ such that $v=P(s,t)$ for a braid relation (1) in $W$, we set $R(s,t):=P(t,s)$.
@@
Let $w_{1},\ldots,w_{r}\in W$ such that $\overline{{B}w_{j}{B}}$ is smooth for all $j=1,\ldots,r$. Suppose for $j\in\{1,\ldots,r\}$, the reduced expression of $w_{j}$ is of the form $w_{j}=P(s,t)$ for a braid relation in $W$. Then for all $i\geq 0$, there are $G^{F}$-equivariant isomorphisms
$H^{i}\left(X_{1},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{2},\mathcal{O}_{X_{2}}\right)$
@@
$H^{i}\left(X_{1},\omega_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{2},\omega_{X_{2}}\right),$
for $X_{n}$ $(n\in\{1,2\})$ being one of the $k$-schemes ${\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$, ${\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$, and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, where $R(s,t)$ replaces $w_{j}$.
@@
By assumption, we have $w_{j}=P(s,t)=R(s,t)$ in $W$, so replacing $w_{j}$ by $R(s,t)$ does not affect the following compactified Deligne–Lusztig variety
$X(\underline{w_{1}},\ldots,\underline{w_{r}})=X(\underline{w_{1}},\ldots,\underline{R(s,t)},\ldots,\underline{w_{r}}).$
Replace $w_{j}$ by $R(s,t)$ in the expression $w_{1}\cdots w_{r}\in S^{*}$ and denote the resulting element by $w_{1}\cdots R(s,t)\cdots w_{r}$. For $X_{n}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$, ${\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, it follows from the assumptions and Lemma 3.1.3 that $X_{n}$ is smooth and projective. Note that we have $I:={\text{supp}}_{F}(w_{1}\cdots w_{r})={\text{supp}}_{F}(w_{1}\cdots R(s,t)\cdots w_{r})={\text{supp}}_{F}(w_{1},\ldots,w_{r})$, and by Proposition 2.3.3 $X_{n}$ has $|G^{F}/P_{I}^{F}|$ irreducible components.
There are natural projection maps ${\rm pr}_{0,\ell(w_{j})}:\overline{O}^{S^{*}}(w_{j})\to O(\underline{w_{j}})$ and ${\rm pr}_{0,\ell(w_{j})}:\overline{O}^{S^{*}}(R(s,t))\to O(\underline{w_{j}})$ for all $j=1,\ldots r$. Similarly, the projection to the 00-th, $\ell(w_{1})$-th, $\ell(w_{1})+\ell(w_{2})$-th, . . ., and $\ell(w_{1})+\cdots+\ell(w_{r})$-th coordinate yields the vertical maps in the cartesian diagrams
${{\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}}$${\overline{O}^{S^{*}}(w_{1}\cdots w_{r})}$${X(\underline{w_{1}},\ldots,\underline{w_{r}})}$${O(\underline{w_{1}},\ldots,\underline{w_{r}}),}$$\scriptstyle{f}$
@@
${{\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}}$${\overline{O}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$${X(\underline{w_{1}},\ldots,\underline{w_{r}})}$${O(\underline{w_{1}},\ldots,\underline{w_{r}}).}$$\scriptstyle{f^{\prime}}$
@@
It follows from the proof of Zariski’s main theorem ([Hartshorne, Corollary 11.4]) that $\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$, for $\phi=f$ (resp. $\phi=f^{\prime}$), $X_{1}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$ (resp. $X_{1}={\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$) and $X_{2}=X(\underline{w_{1}},\ldots,\underline{w_{r}})$. Moreover, applying [CR11, Corollary 3.2.10] yields $R\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$. Hence the Leray spectral sequence associated to $\phi$ degenerates.
@@
for all $i\geq 0$. The $G^{F}$-equivariant isomorphisms for the cohomology of the canonical sheaf follow from an equivariant version of Serre duality [Nironi, Corollary 2.10]. ∎

### v3

@@ 2.3. Criterion for irreducibility
The $F$-orbit of $w\in W$ is the set ${{\rm Orb}_{F}(w)}:=\{F^{n}(w)|n\in\mathbb{Z}_{\geq 0}\}$. We define the set
@@
Let $w_{1},\ldots,w_{r}\in W$ and denote $I={{\text{supp}}}_{F}(w_{1},\ldots,w_{r})$. The Deligne–Lusztig varieties $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are irreducible if and only if $I=S_{F}$. In general, they have $|G^{F}/P_{I}^{F}|$ irreducible components, where $W_{I}$ is the $F$-stable parabolic subgroup of $W$ generated by $\{s\in W|{{\rm Orb}_{F}(s)}\in I\}$ and $P_{I}={B}W_{I}{B}$.
@@
via projection, cf. [8, §2.3.1]. It also follows that $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are dense open subschemes of $\overline{X}^{S^{*}}(v)$}. Thus it suffices to prove that $\overline{X}^{S^{*}}(v)$} is irreducible if and only if $v$ has full $F$-support.
The proof showing that $\overline{X}^{S^{*}}(v)$} is irreducible is analogous to the proof of [3] for classical Deligne–Lusztig varieties. Suppose $I=S_{F}$ and thus ${{\text{supp}}}_{F}(v)=S_{F}$. Let $Z$ be the connected component of $\overline{X}^{S^{*}}(v)$} which contains the point $({B},\ldots,{B})$. Denote the stabiliser of $Z$ in $G^{F}$ by ${\rm Stab}(Z)$. Fix $v^{\prime}$ to be of minimal length such that $v^{\prime}\preceq v$ and ${{\text{supp}}}_{F}(v^{\prime})=S_{F}$. In particular, $v^{\prime}$ is a reduced expression of $\alpha(v^{\prime})\in W$, which is a Coxeter element. This implies that $\overline{X}^{S^{*}}(v^{\prime})$} is irreducible [22, Proposition 4.8].
On the other hand, there is a $G^{F}$-equivariant embedding ${\overline{X}^{S^{*}}(v^{\prime})}\to{\overline{X}^{S^{*}}(v)}$. Let $Z^{\prime}$ be the connected component of $\overline{X}^{S^{*}}(v^{\prime})$} containing $({B},\ldots,{B})$. Then ${\rm Stab}(Z^{\prime})\subseteq{\rm Stab}(Z)$. However, $\overline{X}^{S^{*}}(v^{\prime})$} is irreducible, so ${\rm Stab}(Z^{\prime})=G^{F}$. Thus $\overline{X}^{S^{*}}(v)$} is irreducible. In this case, its dense open subschemes $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are also irreducible [13, Proposition 3.24].
Suppose ${{\text{supp}}}_{F}(w_{1},\ldots,w_{r})=I$ and so ${{\text{supp}}}_{F}(v)=I$. By [8, Proposition 2.3.8], we know that there is a $G^{F}$-equivariant isomorphism
@@
whose fibres are isomorphic to $\overline{X}^{S^{*}}_{L_{I}}(v)$. In this case, we know that $v$ has full $F$-support with respect to the Weyl group of $L_{I}$, so $\overline{X}^{S^{*}}(v)$} has $|G^{F}/P_{I}^{F}|$ irreducible components. ∎
@@ 2.4. $F_{p}$-regularity
In [26, Proposition A.6], we showed that some compactified Deligne–Lusztig varieties are strongly $F_{p}$-regular. This property generalises to arbitrary compactified Deligne–Lusztig varieties. As in loc. cit., we use a similar strategy as in the proof of [23, Lemma 4.3], which uses faithfully flat descent of the smoothness property along the Lang map.
@@
When $\overline{Bw_{i}B}$ is smooth for all $i$, this is proved by faithfully flat descent of smoothness in [23, Lemma 4.3] and [8, Proposition 2.3.5], cf. [13, Proposition 14.59]. Analogously, we use faithfully flat descent of strongly $F_{p}$-regularity to show that any compactified Deligne–Lusztig variety is strongly $F_{p}$-regular.
For each $i=1,\ldots,r$, we know that $\overline{{B}w_{i}{B}}$ is strongly $F_{p}$-regular [4, Cor. 4.1, p. 958] and thus Cohen-Macaulay [4, Cor. 4.2, p. 959]. The product
@@
is strongly $F_{p}$-regular by faithfully flat descent along the projection map ${\rm pr}_{i}:\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}\to\overline{{B}w_{i}{B}}$ [20, Theorem 5.5.b]. Recall that the Lang map $\mathcal{L}_{1}$ is étale, so its closed fibres are smooth and thus strongly $F_{p}$-regular, cf. [25, Ch 3, Remark 1.6, Theorem 4.4(a)]. We may now apply [1, Theorem 3.6] and conclude that
@@
is strongly $F_{p}$-regular. Finally, by Lemma 2.2.1 and the faithfully flat descent of strong $F_{p}$-regularity [20, Theorem 5.5], $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is strongly $F_{p}$-regular.
Strongly $F_{p}$-regular rings are $F_{p}$-rational, so $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is normal and Cohen-Macaulay by [20, Theorem 4.2, 6.27]. ∎
@@
For $w\in W$, the Zariski closure $\overline{X(w)}$ in $G/B$ is not always smooth, but for any $v\in S^{*}$ that gives a reduced expression of $w$, the projection map ${\overline{X}^{S^{*}}(v)}\to\overline{X(w)}$ is a proper birational morphism [7, Lemma 9.11].
@@ 3.1. Smoothness via braid relations
We consider some specific elements in the Weyl group for which the compactified Deligne–Lusztig variety is smooth. To begin, we recall the relations in the Weyl group of finite reductive groups, cf. [11].
All the relations in the group presentation of the Weyl group as a finite Coxeter group $(W,S)$ can be written as $(st)^{m_{st}}=1,s,t\in S,$ where $m_{st}$ is a positive integer such that $m_{ss}=1$ and $m_{st}\geq 2$ if $s\neq t$.
When $(st)^{m}=1$, with $s,t\in S$ and $s\neq t$ is a relation in $W$, we define
$P(s,t):=st\cdots\in S^{*},$
such that ${{\ell^{S^{*}}}(P(s,t))}=m$. When there is no ambiguity, we also use $P(s,t)$ to denote its canonical projection $\alpha(P(s,t))\in W$. Moreover, the equation $(st)^{m}=1$ can be rewritten as
$P(s,t)=P(t,s)$ (1)
with ${{\ell^{S^{*}}}(P(s,t))}={{\ell^{S^{*}}}(P(t,s))}=m$. We call (1) a braid relation for $W$ following the convention of [11, §4].
@@
Let $w_{1},\ldots,w_{r}\in W$. Suppose for each $j\in\{1,\ldots,r\}$, there is a braid relation (1) of $W$ such that $w_{j}=P(s,t)=P(t,s)$. Then the compactified Deligne–Lusztig variety $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth.
@@
By Proposition 2.4.1, it suffices to show that for all $j=1,\ldots,r$, the associated scheme $\overline{{B}w_{j}{B}}$ is smooth.
Consider $w_{j}=P(s,t)=P(t,s)$ given by $(st)^{m}=1$. Denote $J=\{s,t\}\subseteq S$. The the corresponding subgroup $W_{J}\subseteq W$ generated by $J$ has a group presentation
@@
Let $P_{J}=BW_{J}B\subseteq G$ be the parabolic subgroup corresponding to $W_{J}$. Since $w_{j}$ is the longest element in $W_{J}$, we know that $\overline{{B}w_{j}{B}}=P_{J}$ and is smooth cf. [9, Proposition 3.2.8]. ∎
@@ 3.2. Higher direct images of the structure sheaf for birational morphisms
Recall that the braid monoid $B^{+}=B^{+}(W,S)$ associated to $W$ is a monoid generated by $S$ with respect to the braid relations (1) for $W$, cf. [11, §4]. There is a natural surjective morphism of monoids $\beta:S^{*}\twoheadrightarrow B^{+}$. Note that the projection map $\alpha:S^{*}\to W$ from §2 factors through $\beta$.
@@
Let $w,v\in S^{*}$ such that $\beta(w)=\beta(v)$. Then for all $i\geq 0$, there are $G^{F}$-equivariant isomorphisms
$H^{i}\left({\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\right)\overset{\sim}{\longrightarrow}H^{i}\left({\overline{X}^{S^{*}}(v)},\mathcal{O}_{{\overline{X}^{S^{*}}(v)}}\right)$
@@
$H^{i}\left({\overline{X}^{S^{*}}(w)},\omega_{{\overline{X}^{S^{*}}(w)}}\right)\overset{\sim}{\longrightarrow}H^{i}\left({\overline{X}^{S^{*}}(v)},\omega_{{\overline{X}^{S^{*}}(v)}}\right).$
@@
We can rewrite $w=w_{1}\cdots w_{r}$ in $S^{*}$ such that ${{\ell^{S^{*}}}(w_{j})}=\ell(\alpha(w_{j}))$ and $\overline{{B}\alpha(w_{j}){B}}$ is smooth for all $j=1,\ldots,r$. Suppose for an $l\in\{1,\ldots,r\}$, we have $w_{l}=P(s,t)$ given by a braid relation $P(s,t)=P(t,s)$ of $W$. Define the element $w^{\prime}:=w_{1}\cdots P(t,s)\cdots w_{r}$ in $S^{*}$ by replacing $w_{l}$ by $P(t,s)$. We show below that $\overline{X}^{S^{*}}(w)$}, $\overline{X}^{S^{*}}(w^{\prime})$}, and $X(\underline{\alpha(w_{1})},\ldots,\underline{\alpha(w_{r}}))$ have the same cohomology for the structure sheaf and canonical sheaf. Since $\beta(w)=\beta(v)$, the statement of the proposition follows from applying this procedure finitiely many times.
By assumption, we have $\alpha(w_{l})=P(s,t)=P(t,s)$ in $W$, so replacing $w_{l}$ by $P(t,s)$ does not affect the construction of the following compactified Deligne–Lusztig variety
$X(\underline{\alpha(w_{1})},\ldots,\underline{\alpha(w_{r}}))=X(\underline{\alpha(w_{1})},\ldots,\underline{P(t,s)},\ldots,\underline{\alpha(w_{r}})).$
For $X_{n}={\overline{X}^{S^{*}}(w)}$, $\overline{X}^{S^{*}}(w^{\prime})$}, $X(\underline{\alpha(w_{1})},\ldots,\underline{\alpha(w_{r}}))$, it follows from the assumptions and Lemma 3.1.1 that $X_{n}$ is smooth and projective. Note that we have $I:={\text{supp}}_{F}(w)={\text{supp}}_{F}(w^{\prime})={\text{supp}}_{F}(\alpha(w_{1}),\ldots,\alpha(w_{r}))$, so $X_{n}$ has $|G^{F}/P_{I}^{F}|$ irreducible components by Proposition 2.3.3,.
There are natural projection maps ${\rm pr}_{0,\ell(w_{j})}:\overline{O}^{S^{*}}(w_{j})\to O(\underline{w_{j}})$ for all $j=1,\ldots r$ and ${\rm pr}_{0,\ell(w_{l})}:\overline{O}^{S^{*}}(P(t,s))\to O(\underline{\alpha(w_{l})})$. Similarly, the projection to the 0-th, $\ell(w_{1})$-th, $\ell(w_{1})+\ell(w_{2})$-th, . . ., and $\ell(w_{1})+\cdots+\ell(w_{r})$-th coordinate yields the vertical maps in the cartesian diagrams
$\overline{X}^{S^{*}}(w)$$\overline{O}^{S^{*}}(w)$$X(\underline{\alpha(w_{1})},\ldots,\underline{\alpha(w_{r}}))$$O(\underline{\alpha(w_{1})},\ldots,\underline{\alpha(w_{r}})),$$\scriptstyle{f}$
@@
$\overline{X}^{S^{*}}(w^{\prime})$$\overline{O}^{S^{*}}(w^{\prime})$$X(\underline{\alpha(w_{1})},\ldots,\underline{\alpha(w_{r}}))$$O(\underline{\alpha(w_{1})},\ldots,\underline{\alpha(w_{r}})).$$\scriptstyle{f^{\prime}}$
@@
It follows from the proof of Zariski’s main theorem ([17, Corollary 11.4]) that $\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$, for $\phi=f$ (resp. $\phi=f^{\prime}$), $X_{1}={\overline{X}^{S^{*}}(w)}$ (resp. $X_{1}={\overline{X}^{S^{*}}(w^{\prime})}$) and $X_{2}=X(\underline{\alpha(w_{1})},\ldots,\underline{\alpha(w_{r}}))$. Moreover, applying [6, Corollary 3.2.10] yields $R\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$. Hence the Leray spectral sequence associated to $\phi$ degenerates.
@@
for all $i\geq 0$. The $G^{F}$-equivariant isomorphisms for the cohomology of the canonical sheaf follow from an equivariant version of Serre duality [24, Corollary 2.10]. ∎